OpenStudy (anonymous):

If a coin is tossed 3 times, what is the probability of getting 3 tails? A) 1/4 B) 1/9 C) 1/6 D) 1/8

OpenStudy (anonymous):

@apoorvk

OpenStudy (apoorvk):

Hmm, once again - we need a sample space of all the possible outcomes. Can you make one this time? :]

OpenStudy (anonymous):

p(tail) = 1/2 tossed 3 times in a sample space of \[2^{3}\] it becomes: \[p(tail _{3}) = (1/2)^{3} = 1/8\]

OpenStudy (anonymous):

Sample space: TH HT HH TT but three time because of three coins.

OpenStudy (apoorvk):

Umm, actually don't you think that's just for two coins? I can't see outcomes for the third coin in there. Like 'TH' - tail for the first, head for the second, but the third?? ;)

OpenStudy (anonymous):

Oh, haha okay.

OpenStudy (apoorvk):

Hmm, let me show you how: So one easy way to count this is - First I assume Tail for the first coin, and combine it with all the outcomes for '2 coins', then I fix Head, and do the same. Like this: T - HT T - TH T - HH T - TT and H-HT H-TH H-HH H-TT So, basically, I have the following outcomes: THT TTH THH TTT HHT HTH HHH TTT Understandable?

OpenStudy (anonymous):

yes. :3

OpenStudy (apoorvk):

So how many possible outcomes do I have, and how many among those are outcomes with 3 tails?

OpenStudy (apoorvk):

@darkzii ??